The area of a school yard is represented by 8x^2 + 40x. Find the actual dimensions of the school yard if x = 3m

Found 3 solutions by Alan3354, Fombitz, solver91311:
Answer by Alan3354(69443)

The area of a school yard is represented by 8x^2 + 40x. Find the actual dimensions of the school yard if x = 3m
=================
If 3m means 3 meters:
---------
Area = 8x^2 + 40x = 8*9 + 40*3 = 192 sq meters

Answer by Fombitz(32388)
Answer by solver91311(24713)





Factor out where



Substitute



The dimensions of the playground are by where is any positive real number. In other words, there are an infinite number of solutions to this problem.

John

My calculator said it, I believe it, that settles it